What is the partial pressure of sulfur dioxide gas in torr in a sealed vessel that has a total pressure of 520 torr and partial pressures of 0.35 atm for nitrogen?
1
Expert's answer
2017-09-04T14:48:06-0400
P_gen=P(〖SO〗_2 )+P(N_2)
P(〖SO〗_2 )= P_gen-P(N_2)
1 atm = 760 torr
P(N_2 )=(760 torr×0.35 atm)/(1 atm)=266 torr
P(〖SO〗_2 )= 520-266=254 (torr)

Answer: 254 torr.